OpenStudy (anonymous):

How do I put y-2=-1/3(x-1) into slope intercept form?

OpenStudy (amistre64):

simply put, slope intercept form is a format" y = mx + b so you just solve the setup for y in terms of x

OpenStudy (anonymous):

I know that I need to distribute -1/3 but I don't know exactly what to do with -1/3×-1. Would it just end up being y-2=-1/3x+1/3?

OpenStudy (amistre64):

so far, yes

OpenStudy (amistre64):

now add off that 2 to get the y all by itself

OpenStudy (anonymous):

Thank you for your help, I was actually supposed to put it into standard form and got the two mixed up when I wrote it on here. But now I know that there isn't anything fancy with fractions, thanks again.

OpenStudy (amistre64):

:) youre welcome